Use Jevity 1.2 as the enteral TF formula for all cases.

52 y/o Female

Medical Dx: Schizophrenia, CVA, HTN, Depression

Anthropometrics: Current wt:   162 lbs,   ht: 68 in,     UBW: 193# (recent weight loss following CVA)

Activity: bedbound

Physical findings: skin intact, no edema

Biochemical assessment: all pertinent labs WNL.

Diet: NPO

SHOW WORK/FORMULAS

1. Assess the patient’s weight and determine weight goal, including justification.

2. Estimate the patient’s energy, protein, and fluid needs.

3. Calculate the tube feeding.

4. Determine the rate of the continuous feed.

5. How many calories does that feeding provide/day?

6. How many grams protein does this feeding provide/day?

7. How many ml of fluid does this feeding provide/day?

8. Are any additional water flushes needed?

9. Calculate flushes, using 4 flushes/day:

10. Write the final Nutrition Rx for the tube feeding.
Answer #1

1)Given patient weight 162lb(73kg)

Height: 68" (172.7cm)

Expected weight= Height-100

172.7-100=72.7 kg (73kg approximately)

The patient has reached her right goal of weight according to her height

2)The required energy per day can be measured by just multiplying weight by 10.so 73*10=730 calories, but to meet the nutritional needs any female who is bedridden should get a average 1200calories/day

The protein requirement is by multiplying weight with 0.8 73*0.8=58.4 is her requirements

The fluid requirements is generally normal as her blood investigation are WNL 2.2 litre/day

3)The tube feeding should have atleat 10 feeds every second hourly to maintain the nutritional the nutritional status with a calories intake of 1800

4)The rate should be 20-40ml/hour if continuous feed is ordered. In a hour it will be 5- 10 drop/minutes (20*15/1*60=5,40*15/1/60=10)